# Cosmos World Foundation Model Platform for Physical AI

### v1

@@ 6.3.3 Evaluation
| Cosmos-1.0-Diffusion-7B-Text2World- Sample-MultiView | 32.16 | 210.23 | 0.68 | 2.11 |
| Cosmos-1.0-Diffusion-7B-Text2World- Sample-MultiView-TrajectoryCond | - | - | 0.59 | 2.02 |
@@
| Cosmos-1.0-Diffusion-7B-Text2World- Sample-MultiView | 0.77 | 4.25 | 0.29 | - |
| Cosmos-1.0-Diffusion-7B-Text2World- Sample-MultiView-TrajectoryCond | 0.54 | 4.31 | 0.18 | 20.20 |
@@
We first present text-conditioned qualitative results in Fig. 27. Using Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView, we generate a 57-frame video with six views, which is then extended to 201 frames using the Cosmos-1.0-Diffusion-7B-Video2World-Sample-MultiView model. In Fig. 28, we demonstrate how the pre-trained world model enhances generalization, enabling the generation of rare or out-of-domain scenes from the RDS dataset, such as driving on a river. Lastly, Fig. 29 showcases the results from Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView-TrajectoryCond, where the ego car accurately follows the input trajectory.
@@
Generation quality. We utilize Fréchet Inception Distance (FID) (Heusel et al., 2017) and Fréchet Video Distance (FVD) (Unterthiner et al., 2019) to measure the quality of the generated videos relative to the real ones. We first calculate a score per view by extracting 16 frames from each video. We then report the average score across all views per method. As shown in Tab. 24, we find both Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView and Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView-TrajectoryCond significantly outperform VideoLDM-MultiView on both metrics, demonstrating the superior quality of our pre-trained 7B Diffusion-based WFM over the VideoLDM-MultiView baseline.
@@
As shown in Tab. 24, we find that both Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView and Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView-TrajectoryCond render better multi-view geometry consistency over VideoLDM-MultiView. The overall geometric plausibility of the generated videos is much better for a world model fine-tuned from our WFM. We also note that, with the trajectory control condition, such consistency is further improved thanks to the explicit 3D guidance as Cosmos-1.0-Diffusion-7B-Text2World-Sample-MultiView-TrajectoryCond is ranked the best.
@@ 7.2.1 Video Content Safety Filter
We collect three kinds of ground truth annotated data. First, we sample a large set of videos from our dataset, extract frames, and determine its class using a VLM. Next, we generate synthetic videos with our WFMs using a set of prompts to ensure coverage of corner cases and least-represented content categories. Finally, human annotators provide the “gold standard” labels for a portion of our dataset, adding a vital layer of validation and helping us continuously refine the accuracy of our classifier. We extract the SigLIP (Zhai et al., 2023) embedding for each video frame and train a simple MLP classifier on the embeddings.
During inference, we generate a SigLIP embedding for every frame and then apply the classifier. The entire video is flagged as unsafe if any frame is classified as unsafe.
@@ A.3 Acknowledgments
We thank Aarti Basant, Akan Huang, Alex Qi, Alexis Bjorlin, Amanda Moran, Amol Fasale, Ankit Patel, Arash Vahdat, Aryaman Gupta, Ashna Khetan, Ashwath Aithal, Bor-Yiing Su, Bryan Catanzaro, Charles Hsu, Chris Pruett, Christopher Horvath, Clark Doan, Coulten Holt, Dane Aconfora, Deepak Narayanan, Dennis Chang, Dheeraj Kapur, Dong Ahn, Ebrar Erdem, Elmar Haussmann, Gandhi Vaithilingam, Henry Estela, Henry Vera, Herb Woodruff, Imad El Hanafi, Jashojit Mukherjee, Jason Sewall, Jensen Huang, John Dickinson, John Dickinson, Jonah Alben, Jonah Philion, Josh Abbott, Jun Gao, Kumar Anik, Lee Ditiangkin, Luke Alonso, Madison Huang, Marek Dabek, Mark Arnold, Max Ehrlich, Michele Ferretti, Misbah Mubarak, Misha Smelyanskiy, Mohamed Fawzy, Mohammad Harrim, Mohammad Shoeybi, Omkar Mehta, Pallab Bhattacharya, Paniz Karbasi, Pasha Shamis, Raju Wagwani, Rick Izzo, Robert Hero, Sharon Clay, Songyan Tang, Sophia Huang, Sridhar Bhuvanapalli, TJ Galda, Thomas Volk, Tobias Lasser, Vaibhav Ranglani, Vijay Anand Korthikanti, Yazdan Aghaghiri, Yugi Guvvala, and Zekun Hao for their feedback and engineering support.
We thank Iain Cunningham, Jim Fan, Marco Pavone, Meredith Price, Nikki Pope, Scott Reed, and Yuke Zhu for their feedback on the early draft of this technical report.

### v2

@@ 6.3.3 Evaluation
| Cosmos-Predict1-7B-Text2World- Sample-MultiView | 32.16 | 210.23 | 0.68 | 2.11 |
| Cosmos-Predict1-7B-Text2World- Sample-MultiView-TrajectoryCond | - | - | 0.59 | 2.02 |
@@
| Cosmos-Predict1-7B-Text2World- Sample-MultiView | 0.77 | 4.25 | 0.29 | - |
| Cosmos-Predict1-7B-Text2World- Sample-MultiView-TrajectoryCond | 0.54 | 4.31 | 0.18 | 20.20 |
@@
We first present text-conditioned qualitative results in Fig. 27. Using Cosmos-Predict1-7B-Text2World-Sample-MultiView, we generate a 57-frame video with six views, which is then extended to 201 frames using the Cosmos-Predict1-7B-Video2World-Sample-MultiView model. In Fig. 28, we demonstrate how the pre-trained world model enhances generalization, enabling the generation of rare or out-of-domain scenes from the RDS dataset, such as driving on a river. Lastly, Fig. 29 showcases the results from Cosmos-Predict1-7B-Text2World-Sample-MultiView-TrajectoryCond, where the ego car accurately follows the input trajectory.
@@
Generation quality. We utilize Fréchet Inception Distance (FID) (Heusel et al., 2017) and Fréchet Video Distance (FVD) (Unterthiner et al., 2019) to measure the quality of the generated videos relative to the real ones. We first calculate a score per view by extracting 16 frames from each video. We then report the average score across all views per method. As shown in Tab. 24, we find both Cosmos-Predict1-7B-Text2World-Sample-MultiView and Cosmos-Predict1-7B-Text2World-Sample-MultiView-TrajectoryCond significantly outperform VideoLDM-MultiView on both metrics, demonstrating the superior quality of our pre-trained 7B Diffusion-based WFM over the VideoLDM-MultiView baseline.
@@
As shown in Tab. 24, we find that both Cosmos-Predict1-7B-Text2World-Sample-MultiView and Cosmos-Predict1-7B-Text2World-Sample-MultiView-TrajectoryCond render better multi-view geometry consistency over VideoLDM-MultiView. The overall geometric plausibility of the generated videos is much better for a world model fine-tuned from our WFM. We also note that, with the trajectory control condition, such consistency is further improved thanks to the explicit 3D guidance as Cosmos-Predict1-7B-Text2World-Sample-MultiView-TrajectoryCond is ranked the best.
@@ 7.2.1 Video Content Safety Filter
We collect three kinds of ground truth annotated data. First, we sample a large set of videos from our dataset, extract frames, and determine its class using a VLM. Next, we generate synthetic videos with our WFMs using a set of prompts to ensure coverage of corner cases and least-represented content categories. Finally, human annotators provide the “gold standard” labels for a portion of our dataset, adding a vital layer of validation and helping us continuously refine the accuracy of our classifier. We extract the SigLIP (Zhai et al., 2023) embedding for each video frame and train a simple MLP classifier on the embeddings. ing inference, we generate a SigLIP embedding for every frame and then apply the classifier. The entire video is flagged as unsafe if any frame is classified as unsafe
@@ A.3 Acknowledgments
We thank Aarti Basant, Akan Huang, Alex Qi, Alexis Bjorlin, Amanda Moran, Amol Fasale, Ankit Patel, Arash Vahdat, Aryaman Gupta, Ashna Khetan, Ashwath Aithal, Bor-Yiing Su, Bryan Catanzaro, Charles Hsu, Chris Pruett, Christopher Horvath, Clark Doan, Coulten Holt, Dane Aconfora, Deepak Narayanan, Dennis Chang, Dheeraj Kapur, Dong Ahn, Ebrar Erdem, Elmar Haussmann, Fuzhao Xue, Gandhi Vaithilingam, Henry Estela, Henry Vera, Herb Woodruff, Imad El Hanafi, Jashojit Mukherjee, Jason Sewall, Jensen Huang, John Dickinson, Jonah Alben, Jonah Philion, Josh Abbott, Jun Gao, Kumar Anik, Lee Ditiangkin, Ligeng Zhu, Linxi Fan, Luke Alonso, Madison Huang, Marek Dabek, Mark Arnold, Max Ehrlich, Michele Ferretti, Misbah Mubarak, Misha Smelyanskiy, Mohamed Fawzy, Mohammad Harrim, Mohammad Shoeybi, Omkar Mehta, Pallab Bhattacharya, Paniz Karbasi, Pasha Shamis, Raju Wagwani, Rick Izzo, Robert Hero, Sharon Clay, Song Han, Songyan Tang, Sophia Huang, Sridhar Bhuvanapalli, TJ Galda, Thomas Volk, Tobias Lasser, Vaibhav Ranglani, Vijay Anand Korthikanti, Yao Lu, Yazdan Aghaghiri, Yugi Guvvala, Yuke Zhu and Zekun Hao for their feedback and engineering support.
We thank Iain Cunningham, Jim Fan, Marco Pavone, Meredith Price, Nikki Pope and Scott Reed for their feedback on the early draft of this technical report.